# Latent Optimal Paths by Gumbel Propagation for Variational Bayesian Dynamic Programming

### v2

@@ Appendix A Proofs of each lemma in Gumbel propagation
where Equation 40 follows from the identity
@@ Appendix D Gumbel Softmax Trick
In this section, we investigates a possible interpretation of achieving reparameterization trick (i.e., Gumbel softmax trick) on BDP-VAE framework for interested readers. Compared to the log-derivative trick discussed in Section 5.3, Gumbel softmax trick samples soft latent optimal paths and has advantages on gradient-based optimization. However, this method may requires extra effort in designing its implementation and temperature paramter $\tau$.
@@ D.1 Node-wise Gumbel Softmax Propagation
Due to the generally intractable size $\left|\mathcal{Y}(1,N)\right|$ of the set of paths that make up the domain of $\mathcal{D}(\mathbf{y}|\mathcal{R},\mathbf{W},\alpha)$, there is no simple analogue of the above for our setting. Instead, we offer two alternative approaches, both of which are differentiable reparameterisations of a distribution of real values, one per node.
@@ D.1.1 Marginal Node-wise Gumbel Softmax Distribution
Let $\mathbf{y}\sim\mathcal{D}(\mathcal{R},\mathbf{W},\alpha)$. The hitting probability for node $v$ is the probability that $\mathbf{y}$ includes v,
@@ E.1 Dynamic Time Warping
We first extend the Bayesian dynamic programming on the dynamic time warping (DTW) algorithm (Sakoe & Chiba, 1978; Mensch & Blondel, 2018) which aims to seek an optimal alignment path with the maximum score (or minimum cost) given two time series. Given two time-series $\mathbf{A}$ and $\mathbf{B}$ with lengths $N_{A}$ and $N_{B}$. Let $a_{j}$ and $b_{i}$ be the $j^{th}$ and $i^{th}$ observations of $\mathbf{A}$ and $\mathbf{B}$, respectively. We denote $\mathbf{W}\in\mathbb{R}^{N_{B}\times N_{A}}$ as a pair-wise weight matrix, where $w_{ij}$ represents similarity measurement between observation point $b_{i}$ and observation point $a_{j}$, namely, $w_{ij}=d(b_{i},a_{j})$, where function $d(\cdot)$ is an arbitrary similarity metric. The $\mathcal{Y}$ defines the set of the population of all possible time-series alignment paths $\mathbf{y}$, in which the path connects the upper-left $(1,1)$ node to the lower-right $(N_{B},N_{A})$ node with $\rightarrow$, $\downarrow$ and $\searrow$ moves only.
@@
     $\pi_{i,j}=[\frac{\exp(\mu_{i,j-1}+\alpha w_{i,j})}{\exp(\mu_{i,j})},\frac{\exp(\mu_{i-1,j-1}+\alpha w_{i,j})}{\exp(\mu_{i,j})},$
     $\frac{\exp(\mu_{i-1,j}+\alpha w_{i,j})}{\exp(\mu_{i,j})}]$
@@
Algorithm 2 Sample alignment path $y$ on DTW
@@
We denote the marginal probability of edges between node $(u,u^{\prime})$ and node $(v,v^{\prime})$, where $(u,u^{\prime})\in\mathcal{P}(v,v)$, as $\omega_{(u,u^{\prime})(v,v^{\prime})}$. Following Lemma 4.9, the $\omega$ can be computed by Equation 86
@@ E.1.1 Time complexity analysis
Following above definition of the computational graph of DTW, where the pair-wise weight matrix $\mathbf{W}\in\mathbb{R}^{N_{B}\times N_{A}}$. According to Corollary 4.11, the edge numbers under DTW graph is $|\mathcal{E}|=3N_{A}N_{B}-2N_{A}-2N_{B}+1$. Therefore, the time complexity in above pseudo algorithm is $\mathcal{O}(3N_{A}N_{B}-2N_{A}-2N_{B}+1)$. Inspired by Tralie & Dempsey (2020), we further reduce its time complexity by updating diagonally in parallel as illustrated in Figure 7. In this way, the time complexity of computational graph of DTW becomes linear $N_{B}$ and $N_{A}$ as $\mathcal{O}(N_{A}+N_{B}-1)$.
@@ E.2 Monotonic Alignment
Monotonic alignment (MA) is often used in the field of machine learning, particularly in the context of sequence-to-sequence (Seq2Seq) models. Taking the same definition of two given time-series in Section E.1, $\mathcal{Y}$ defines a population of all possible path sample $\mathbf{y}$, in which the path connects the upper-left $(1,1)$ node to the lower-right $(N_{B},N_{A})$ node with $\rightarrow$ and $\searrow$ moves only, where $N_{B}<N_{A}$.
The location parameters $\mu\in\mathbb{R}^{N_{B}\times N_{A}}$ can be computed following the Lemma 4.4 and the optimal path $\mathbf{y}$ can be sampled according to the transition matrix $\pi\in\mathbb{R}^{N_{B}\times N_{A}\times 2}$. The probability of transition $(v,v^{\prime})\rightarrow(u,u^{\prime})$ for $(u,u^{\prime})\in\mathcal{P}(v,v^{\prime})$ is defined in Equation 85. Figure 8 gives an example of the monotonic alignment computational graph with $N_{A}=7$ and $N_{B}=4$. The bold black arrows indicate one possible aligned path.
@@
Algorithm 5 Sample alignment path $y$ on MA
@@
Pseudocode of computing the location parameter $\mu$ and transition matrix $\pi$ are provided in Algorithm 4. The sampling algorithm is presented in Algorithm 5.
@@
We then denote the marginal probability of edges between node $(u,u^{\prime})$ and node $(v,v^{\prime})$, where $(u,u^{\prime})\in\mathcal{P}(v,v^{\prime})$, as $\omega_{(u,u^{\prime}),(v,v^{\prime})}$. The $\omega$ can be computed by Equation 86 and the pseudocode to compute the marginal probabilities $\omega$ show on Algorithm 6.
@@ E.2.1 Time complexity analysis
Following the setting of computational graph of MA in above, where the pair-wise weight matrix $\mathbf{W}\in\mathbb{R}^{N_{B}\times N_{A}}$ and the graph is defined in Figure 8, where $N_{B}<N_{A}$. The edge number under MA graph is $|\mathcal{E}|=N_{B}N_{A}-N_{B}^{2}$ and the time complexity in pseudo algorithms of computational graph of MA is $\mathcal{O}(N_{B}N_{A}-N_{B}^{2})$. By updating the graph vertically as illustrated in Figure 9, the time complexity can be further reduced to $\mathcal{O}(N_{A}-1)$.
@@ Appendix F Details of model architecture in experiments
Figure 10: Architecture of BDP-VAE on experiments. The red lines are only turned on during training. Black lines stand for the training process while yellow dotted lines stand for the inference phase. The scissors represent the gradient not being back-propagated along with the arrow.
Given a spectrogram input $\mathbf{x}=[x_{1},...,x_{t}]$ and a corresponding phoneme text $\mathbf{c}^{\prime}=[c^{\prime}_{1},...,c^{\prime}_{n}]$ , where $t$ and $n$ are the lengths of the input sequences. We assume that there is an unobserved alignment representation $\mathbf{y}\in\mathbb{R}^{t\times n}$ aligns $\mathbf{x}$ and $\mathbf{c}$ by $\{0,1\}$ only.
@@
The text encoder is used to extract a higher level of linguistic feature $\mathbf{c}\in\mathbb{R}^{n\times 256}$ from phoneme text, which adopts the same structures as the one in FastSpeech2 (Ren et al., 2020), which contains 4 Feed-forward Transformer blocks with 2 multi-head attentions.
@@
The architecture of the decoder is also the same as the one in FastSpeech2 (Ren et al., 2020). In the decoder (Appendix F), the alignment latent $\mathbf{y}$ and the linguistic feature $\mathbf{c}$ are extended by a matrix multiplication which extends the linguistic feature from length $n$ to length $t$ according to the alignment information, then followed by 4 Feed-forward Transformer blocks with 2 multi-head attentions. We add a residual connection after the linear layer with a PostNet to get the final output.
@@
The prior encoder $p(\mathbf{y}|\mathbf{c};\theta)=\mathcal{D}(\mathbf{y}|\mathcal{R},\mathbf{W}^{(0)}=d(f_{\theta}(\mathbf{c}),\mathbf{c}),\alpha)$. Following Lu et al. (2021); Ma et al. (2019), we make use of a Glow (Kingma & Dhariwal, 2018) structure to infer spectrogram features condition on linguistic features. It consists of multiple Glow blocks. Each of the blocks has an actnorm layer, an invertible $1\times 1$ convolutional layer, and an affine-coupling layer. The transformation network in the affine-coupling layer is based on the Transformer decoder which the spectrogram feature $\text{NN}_{\phi}(\mathbf{x})$ as the query and the linguistic feature $\mathbf{c}$ as the key and value. During training, we use the backward pass to infer the probability of the KL divergence. The forward pass is used to generate spectrogram features from the condition and then form the weight matrix to sample latent $\mathbf{\hat{y}}$ during inference. Details of the architecture are in Appendix F.
@@
Following Lu et al. (2021), the length predictor consists of a 1-channel fully connected layer with ReLU activation. The length predictor is optimized by an MSE loss and the gradient will not propagate to the text encoder.
@@ G.2 Experimental Details for End-to-end Text-to-speech
In the evaluation part, we obtain the DTW-MCD score by the package of pymcd.mcd [^3] https://github.com/chenqi008/V2C. The 60-iteration Grinffin-Lim Algorithm approximates all the synthesized waveforms. We trained all the models in Table 1 with the same pre-processing setting and the details are:
@@
VAENAR-TTS VAENAR-TTS (Lu et al., 2021) is an end-to-end utterance-level TTS model that relies on cascade attention in the decoder of the VAE to obtain the phoneme duration alignment. We followed the model configuration the paper provided and trained the model by 150k iterations and the total loss converged at around $7.4\text{e}^{-3}$.
@@ Appendix H Interpretation with the comparison of VAENAR-TTS
Both BDPVAE-TTS and VAENAR-TTS predict phone-duration alignment on the utterance level. VAENAR-TTS learns the Gaussian latent distribution of learned features of spectrograms and conditions and obtains soft monotonic phoneme-utterance alignment by attention with a causality mask in VAE’s transformer decoder. Different from VAENAR-TTS, BDPVAE-TTS was adapted from a non-end-to-end phoneme-level TTS model (FastSpeech2 (Ren et al., 2020)) with a distribution of stochastic optimal paths $\mathcal{D}(\mathcal{R},\cdot,\alpha)$ into a BDP-VAE framework. Thus, during its inference time, BDPVAE-TTS reconstructs the spectrogram according to phoneme duration aligned phoneme tokens only. Since our baseline method (i.e., FastSpeech2) did not outperform the VAENAR-TTS, thus, BDPVAE-TTS also gets a lower performance than VAENAR-TTS.
However, soft alignment in VAENAR-TTS may make the model cannot accurately capture the exact relationship of how phoneme tokens perform in a spectrogram. According to the natural inherent relationship of speech and phonemes, the BDPVAE-TTS obtains sparse monotonical optimal paths in the latent space which are more precise in explaining the relationship between utterance and phoneme tokens. Figure 13 to Figure 13 give additional demonstrations that the synthesized audios from BDPVAE-TTS have closer F0 to the ground truth than VAENAR-TTS.

### v3

@@ Appendix A Proofs of each lemma in Gumbel propagation
where Equation 36 restates (15), Equation 37 follows from (35), Equation 38 expands the summation, Equation 39 follows from the definition of $\left\|\mathbf{y}\right\|_{\mathbf{W}}$, Equation 40 follows from the identity
@@ Appendix D Gumbel Softmax Trick
In this section, we investigate a possible interpretation of achieving reparameterization trick (i.e., Gumbel softmax trick) on BDP-VAE framework for interested readers. Compared to the log-derivative trick discussed in Section 5.3, Gumbel softmax trick samples soft latent optimal paths and has advantages on gradient-based optimization. However, this method may require extra effort in designing its implementation and temperature parameter $\tau$.
@@ D.1 Node-wise Gumbel Softmax Propagation
Due to the generally intractable size $\left|\mathcal{Y}(1,N)\right|$ of the set of paths that make up the domain of $\mathcal{D}(\mathbf{y}|\mathcal{R},\mathbf{W},\alpha)$, there is no simple analogue of the above for our setting. Instead, we offer two alternative approaches, both of which are differentiable reparameterizations of a distribution of real values, one per node.
@@ D.1.1 Marginal Node-wise Gumbel Softmax Distribution
Let $\mathbf{y}\sim\mathcal{D}(\mathcal{R},\mathbf{W},\alpha)$. The hitting probability for the node $v$ is the probability that $\mathbf{y}$ includes v,
@@ E.1 Dynamic Time Warping
We first extend the Bayesian dynamic programming to the dynamic time warping (DTW) algorithm (Sakoe & Chiba, 1978; Mensch & Blondel, 2018) which aims to seek an optimal alignment path with the maximum score (or minimum cost) given two time series. Given two time-series $\mathbf{A}$ and $\mathbf{B}$ with lengths $N_{A}$ and $N_{B}$. Let $a_{j}$ and $b_{i}$ be $j^{th}$ and $i^{th}$ observations of $\mathbf{A}$ and $\mathbf{B}$, respectively. We denote $\mathbf{W}\in\mathbb{R}^{N_{B}\times N_{A}}$ as a pair-wise weight matrix, where $w_{ij}$ represents similarity measurement between observation point $b_{i}$ and observation point $a_{j}$, namely, $w_{ij}=d(b_{i},a_{j})$, where function $d(\cdot)$ is an arbitrary similarity metric. The $\mathcal{Y}$ defines the set of the population of all possible time-series alignment paths $\mathbf{y}$, in which the path connects the upper-left $(1,1)$ node to the lower-right $(N_{B},N_{A})$ node with $\rightarrow$, $\downarrow$ and $\searrow$ moves only.
@@
     $\pi_{i,j}=[\frac{\exp(\mu_{i,j-1}+\alpha w_{i,j})}{\exp(\mu_{i,j})},\frac{\exp(\mu_{i-1,j-1}+\alpha w_{i,j})}{\exp(\mu_{i,j})},\frac{\exp(\mu_{i-1,j}+\alpha w_{i,j})}{\exp(\mu_{i,j})}]$
@@
Algorithm 2 Sample stochastic optimal path $y$ on DTW
@@
We denote the marginal probability of edges between node $(u,u^{\prime})$ and node $(v,v^{\prime})$, where $(u,u^{\prime})\in\mathcal{P}(v,v)$, as $\omega_{(u,u^{\prime})(v,v^{\prime})}$. Following Lemma 4.9, $\omega$ can be computed by Equation 86
@@ E.1.1 Time complexity analysis
Following above definition of the computational graph of DTW, where the pair-wise weight matrix $\mathbf{W}\in\mathbb{R}^{N_{B}\times N_{A}}$. According to Corollary 4.11, the edge numbers under DTW graph is $|\mathcal{E}|=3N_{A}N_{B}-2N_{A}-2N_{B}+1$. Therefore, the time complexity in the above pseudo algorithm is $\mathcal{O}(3N_{A}N_{B}-2N_{A}-2N_{B}+1)$. Inspired by Tralie & Dempsey (2020), we further reduce its time complexity by updating diagonally in parallel as illustrated in Figure 7. In this way, the time complexity of computational graph of DTW becomes linear $N_{B}$ and $N_{A}$ as $\mathcal{O}(N_{A}+N_{B}-1)$.
@@ E.2 Monotonic Alignment
Monotonic alignment (MA) is often used in the field of machine learning, particularly in the context of sequence-to-sequence (Seq2Seq) tasks. Taking the same definition of two given time-series in Section E.1, $\mathcal{Y}$ defines a population of all possible path sample $\mathbf{y}$, in which the path connects the upper-left $(1,1)$ node to the lower-right $(N_{B},N_{A})$ node with $\rightarrow$ and $\searrow$ moves only, where $N_{B}<N_{A}$.
The location parameters $\mu\in\mathbb{R}^{N_{B}\times N_{A}}$ can be computed following Lemma 4.4 and the optimal path $\mathbf{y}$ can be sampled according to the transition matrix $\pi\in\mathbb{R}^{N_{B}\times N_{A}\times 2}$. The probability of transition $(v,v^{\prime})\rightarrow(u,u^{\prime})$ for $(u,u^{\prime})\in\mathcal{P}(v,v^{\prime})$ is defined in Equation 85. Figure 8 gives an example of the monotonic alignment computational graph with $N_{A}=7$ and $N_{B}=4$. The bold black arrows indicate one possible aligned path.
@@
Algorithm 5 Sample stochastic optimal path $y$ on MA
@@
Pseudo-code for computing the location parameter $\mu$ and transition matrix $\pi$ are provided in Algorithm 4. The sampling algorithm is presented in Algorithm 5.
@@
We then denote the marginal probability of edges between node $(u,u^{\prime})$ and node $(v,v^{\prime})$, where $(u,u^{\prime})\in\mathcal{P}(v,v^{\prime})$, as $\omega_{(u,u^{\prime}),(v,v^{\prime})}$. $\omega$ can be computed by Equation 86 and the pseudocode to compute the marginal probabilities $\omega$ show on Algorithm 6.
@@ E.2.1 Time complexity analysis
Following the setting of computational graph of MA in above, where the pair-wise weight matrix $\mathbf{W}\in\mathbb{R}^{N_{B}\times N_{A}}$ and the graph is defined in Figure 8, where $N_{B}<N_{A}$. The edge number under MA graph is $|\mathcal{E}|=2N_{B}N_{A}-2N_{B}^{2}-N_{A}+2N_{B}-1$ and the time complexity in pseudo algorithms of computational graph of MA is $\mathcal{O}(2N_{B}N_{A}-2N_{B}^{2}-N_{A}+2N_{B}-1)$. By updating the graph vertically as illustrated in Figure 9, the time complexity can be further reduced to $\mathcal{O}(N_{A}-1)$.
@@ Appendix F Details of model architecture in experiments
Figure 10: Architecture of BDP-VAE on experiments. The red lines are only turned on during training. Black lines stand for the training process, while yellow dotted lines stand for the inference phase. The scissors represent the gradient not being back-propagated along with the arrow.
Given a spectrogram input $\mathbf{x}=[x_{1},...,x_{t}]$ and a corresponding phoneme text $\mathbf{c}^{\prime}=[c^{\prime}_{1},...,c^{\prime}_{n}]$ , where $t$ and $n$ are the lengths of the input sequences. We assume that there is an unobserved structural sparse optimal path (i.e., monotonic alignment) representation $\mathbf{y}\in\mathbb{R}^{t\times n}$ aligns $\mathbf{x}$ and $\mathbf{c}$ by $\{0,1\}$ under defined DAG $\mathcal{R}$.
@@
The text encoder is used to extract a higher level of linguistic features $\mathbf{c}\in\mathbb{R}^{n\times 256}$ from phoneme text, which adopts the same structures as the one in FastSpeech2 (Ren et al., 2020), which contains 4 feed-forward transformer (FFT) blocks with 2 multi-head attentions.
@@
The architecture of the decoder is also the same as the one in FastSpeech2 (Ren et al., 2020). In the decoder (Appendix F), the latent optimal path $\mathbf{y}$ and the linguistic feature $\mathbf{c}$ are extended by a matrix multiplication which extends the linguistic feature from length $n$ to length $t$ according to the information of sampled latent optimal path $\mathbf{y}$, then followed by 4 Feed-forward Transformer blocks with 2 multi-head attentions. We add a residual connection after the linear layer with a PostNet to get the final output.
@@
The prior encoder $p(\mathbf{y}|\mathbf{c};\theta)=\mathcal{D}(\mathbf{y}|\mathcal{R},\mathbf{W}^{(0)}=d(f_{\theta}(\mathbf{c}),\mathbf{c}),\alpha)$. Following Lu et al. (2021); Ma et al. (2019), we make use of a Glow (Kingma & Dhariwal, 2018) structure to infer spectrogram features conditioned on linguistic features. It consists of multiple Glow blocks. Each of the blocks has an actnorm layer, an invertible $1\times 1$ convolutional layer, and an affine-coupling layer. The transformation network in the affine-coupling layer is based on the Transformer decoder, which the spectrogram feature $\text{NN}_{\phi}(\mathbf{x})$ as the query and the linguistic feature $\mathbf{c}$ as the key and value. During training, we use the backward pass to infer the probability of the KL divergence. The forward pass is used to generate spectrogram features from the condition and then form the weight matrix to sample latent $\mathbf{\hat{y}}$ during inference. Details of the architecture are in Appendix F.
@@
Following Lu et al. (2021), the length predictor consists of a 1-channel fully connected layer with ReLU activation. The length predictor is optimized by an MSE loss, and the gradient will not propagate to the text encoder.
@@ G.2 Experimental Details for End-to-end Text-to-speech
In the evaluation part, we obtain the DTW-MCD score by the package of pymcd.mcd [^4] https://github.com/chenqi008/V2C. The 60-iteration Grinffin-Lim Algorithm approximates all the synthesized waveforms. We trained all the models in Table 1 with the same pre-processing setting, and the details are:
@@
VAENAR-TTS VAENAR-TTS (Lu et al., 2021) is an end-to-end utterance-level TTS model that relies on cascade-masked self-attention in the decoder of the VAE to obtain the phoneme duration alignment. We followed the model configuration the paper provided and trained the model by 150k iterations and the total loss converged at around $7.4\text{e}^{-3}$.
@@ Appendix H Interpretation with the comparison of VAENAR-TTS
Both BDPVAE-TTS and VAENAR-TTS predict phone-duration alignment on the utterance level. VAENAR-TTS learns the Gaussian latent distribution of learned features of spectrograms and conditions and obtains soft monotonic phoneme-utterance alignment by self-attention with a causality mask in the transformer decoder. Different from VAENAR-TTS, BDPVAE-TTS was adapted from a non-end-to-end phoneme-level TTS model (FastSpeech2 (Ren et al., 2020)) with a Gibbs distribution of stochastic optimal paths $\mathcal{D}(\mathcal{R},\cdot,\alpha)$ defined in Definition 4.1 under monotonic alignment DAG defined in Section E.2 into a BDP-VAE framework. In the inference phase, BDPVAE-TTS reconstructs the spectrogram according to duration-extended phoneme sequences only. Since our baseline method (i.e., FastSpeech2) did not outperform the VAENAR-TTS, thus, BDPVAE-TTS also gets a lower performance than VAENAR-TTS.
However, soft monotonic alignment in VAENAR-TTS may make the model cannot accurately capture the exact relationship of how phoneme tokens perform in a spectrogram. According to the natural inherent relationship of speech and phonemes, the BDPVAE-TTS obtains sparse monotonic optimal paths in the latent space which are more precise in explaining the relationship between utterance and phoneme tokens. Figure 13 to Figure 13 give additional demonstrations that the synthesized audios from BDPVAE-TTS have closer F0 to the ground truth than VAENAR-TTS.